1 - Find an equation of the vertical line that passes through (x, y) = (−10, 18).

2 - Find an equation of the line that passes through the point (−1, 6) and is parallel to the line passing through the points (−3, −5) and (1, 3). (Let x be the independent variable and y be the dependent variable.)

3 - Find an equation of the line that passes through the point (4, 3) and is perpendicular to the line 8x + 5y − 7 = 0. (Let x be the independent variable and y be the dependent variable.)

4 - Using data compiled by the Admissions Office at an imaginary university, college admissions officers estimate that 65% of the students who are offered admissions to the freshman class at the university will actually enroll.

(4a) Find an equation that expresses the relationship between the number of students who actually enroll (y) and the number of students who are offered admission to the university (x).

(4b) If the desired freshman class size for the upcoming academic year is 1300 students, how many students should be admitted?
Answer #1

v vertical line through (x,y) = (-10, 18) - is line i vertical. slope is m= 1 (Not defined) .. y - y, = y- 18 = m(x-x) (x^(-1